OpenStudy (anonymous):

Find the x-intercept(s) and vertex for the quadratic function. f(x)=3x^2-2x-3 Round to the nearest hundredth if necessary. If there is more than one x-intercept, separate them with commas

OpenStudy (luigi0210):

x-intercept is when y=0 \[\LARGE 0=3x^2-2x-3\] Factor and solve or use the quadratic formula.

OpenStudy (zzr0ck3r):

vertex for the trinomial \(f(x) = ax^2+bx+c\) is given by \(V= (\frac{-b}{2a},f(\frac{-b}{2a}))\)

OpenStudy (luigi0210):

Oh, yea vertex formula too^

OpenStudy (anonymous):

so, x intercept= 1.39, 0.72?

OpenStudy (zzr0ck3r):

I would leave it as the exact answer (fraction).

OpenStudy (anonymous):

even though it tells me to round to the nearest hundredth?

OpenStudy (zzr0ck3r):

oh disregard that comment then

OpenStudy (anonymous):

0.33 for the vertex? @Luigi0210

OpenStudy (anonymous):

0.33, -3.33 for the vertex